### v1

# VPTracker: Global Vision-Language Tracking via Visual Prompt and MLLM

Abstract
Vision-Language Tracking aims to continuously localize objects described by a visual template and a language description. Existing methods, however, are typically limited to local search, making them prone to failures under viewpoint changes, occlusions, and rapid target movements. In this work, we introduce the first global tracking framework based on Multimodal Large Language Models (VPTracker), exploiting their powerful semantic reasoning to locate targets across the entire image space. While global search improves robustness and reduces drift, it also introduces distractions from visually or semantically similar objects. To address this, we propose a location-aware visual prompting mechanism that incorporates spatial priors into the MLLM. Specifically, we construct a region-level prompt based on the target’s previous location, enabling the model to prioritize region-level recognition and resort to global inference only when necessary. This design retains the advantages of global tracking while effectively suppressing interference from distracting visual content. Extensive experiments show that our approach significantly enhances tracking stability and target disambiguation under challenging scenarios, opening a new avenue for integrating MLLMs into visual tracking. Code is available at https://github.com/jcwang0602/VPTracker.
I Introduction
Visual object tracking (VOT) is a fundamental problem in computer vision, supporting a wide range of downstream applications such as intelligent surveillance, human–robot interaction, augmented reality, and autonomous systems [10149530]. Traditional visual trackers primarily rely on an initial visual template and perform frame-by-frame localization using appearance similarity [ye2022ostrack]. Although such methods have achieved strong short-term performance, they are inherently constrained by the limited representational capacity of a single visual template. As a result, their robustness deteriorates in complex scenarios involving drastic appearance changes, heavy occlusions, large camera motion, or long-term absence.
To overcome the ambiguity of purely visual cues, vision-language tracking has recently emerged as a promising direction [11209477]. By incorporating natural language descriptions, trackers can leverage high-level semantic priors to better distinguish objects with similar appearance and maintain identity consistency over time. This paradigm has shown particular advantages in multi-object scenes and cases where visual cues alone are insufficient [wang2021towards]. However, existing most vision-language trackers still predominantly adopt local search strategies, restricting the search region to a small neighborhood around the previous target location. While efficient, this localized assumption makes the tracker vulnerable to drift: once the target moves rapidly, undergoes abrupt viewpoint changes, or becomes temporarily occluded, the tracker may fail to rediscover it.
Recent advancements in Multimodal Large Language Models (MLLMs) have opened new possibilities for rethinking the tracking pipeline. MLLMs exhibit remarkable capabilities in global scene understanding, cross-modal reasoning [bai2025qwen3]. These strengths suggest that MLLMs could fundamentally reshape long-term tracking, allowing the model to perform global search over the entire image rather than relying on local heuristics. In this paper, we present the first global vision–language tracking framework powered by Multimodal Large Language Models (MLLMs), termed VPTracker. Leveraging the strong semantic reasoning capabilities of MLLMs, our approach can localize the target across the entire image space. Despite their impressive semantic understanding, we observe that large-scale language models inherently lack temporal modeling and spatial continuity awareness—two key elements required for reliable visual tracking. To overcome this limitation, we introduce a location-aware visual prompting mechanism that injects spatial priors directly into the MLLM’s inference process. Specifically, we construct a region-level prompt based on the target’s previous location, guiding the model to focus on the most probable search area while still maintaining the ability to perform global reasoning when necessary. This design preserves the robustness of global search while substantially reducing false positives caused by visually or semantically similar distractors.
The experimental results on the commonly used visual language tracking datasets TNL2K [wang2021towards], and TNLLT [wang2025reasoningtrack] demonstrate that our method has achieved significant improvements in robustness, target re-identification, and long-term stability. More importantly, VPTracker introduces a new perspective of integrating MLLM into vision-language tracking tasks, indicating that future tracking will not only be driven by appearance similarity but also by advanced cross-modal reasoning. In summary, the main contributions of this work are threefold:
  • We propose the first global vision–language tracking framework built on Multimodal Large Language Models (MLLMs), enabling global target localization across the entire image space rather than relying on traditional local search heuristics. This introduces a new paradigm for long-term tracking driven by cross-modal semantic reasoning.
  • We identify the lack of temporal modeling and spatial continuity in MLLMs and address it with a location-aware visual prompting mechanism. By injecting region-level spatial priors derived from the target’s previous state, our method significantly reduces false positives and enhances target recognition while preserving global reasoning capabilities.
  • Extensive experiments on TNL2K and TNLLT demonstrate that our approach achieves superior robustness, improved recognition capability, and strong long-term stability, establishing a new direction for integrating MLLMs into vision–language tracking..
II Related Work
II-A Object Tracking
Visual object tracking (VOT) has been extensively studied over the past decade, with research evolving from correlation filter–based methods to modern deep learning–based frameworks [ye2022ostrack]. Early trackers such as KCF [henriques2014high] and ECO [danelljan2017eco] model target appearance through handcrafted features and efficient frequency-domain operations. With the success of deep convolutional networks, Siamese-based trackers (e.g., SiamFC [bertinetto2016fully], SiamRPN [li2018high], SiamMask [hu2023siammask]) became dominant due to their strong template-matching capability and real-time performance. Transformer-based trackers (e.g., TransT [chen2021TransT], STARK [tang2022SFTransT], OSTrack [ye2022ostrack]) further improve robustness by modeling long-range dependencies with self-attention mechanisms.
The introduction of natural language into visual tracking provides richer semantic information beyond appearance cues. Early works such as MANet and ReferFormer incorporate text embeddings into the tracking pipeline to better distinguish visually similar objects. LST, TNL2K-trackers, and subsequent models integrate joint vision-language encoders (e.g., CLIP) to enhance target localization through semantic alignment. These approaches demonstrate the advantages of textual descriptions in improving identity consistency.
However, whether in visual tracking or vision–language tracking, most existing methods still follow a local search paradigm, where the search region is restricted to a fixed-size window centered around the last known target position. When the target undergoes large displacement or disappears due to occlusion, the tracker often fails to re-detect it. Moreover, most existing methods rely on shallow fusion of linguistic and visual features, restricting the model’s ability to reason globally about object semantics, inter-object relationships, and scene-level context. As a result, they struggle in complex environments with distractors that share similar descriptions or visual patterns.
II-B Multimodal Large Language Models
Multimodal Large Language Models (MLLMs), such as GPT-4V [achiam2023gpt], LLaVA [li2024llava], Qwen-VL[bai2025qwen3], and InternVL [wang2025internvl3], have exhibited strong capabilities in high-level scene understanding, cross-modal reasoning, and open-vocabulary recognition [wang2025unlocking]. Unlike traditional vision-language models that output fixed embeddings, MLLMs are able to perform explicit reasoning, generate spatial descriptions, and interpret complex relationships between objects. These properties make them promising for challenging video tasks such as grounding, captioning, and instruction-following.
Nevertheless, applying MLLMs directly to visual tracking remains underexplored. Existing works primarily focus on isolated tasks such as image-level grounding or referring expression comprehension, without considering temporal consistency, spatial priors, or the recursive nature of tracking. Moreover, MLLMs often favor global but semantically-driven interpretation, which can lead to ambiguous predictions when visually similar objects appear in the scene. This reveals a fundamental gap between the global semantic reasoning strengths of MLLMs and the precise, temporally grounded localization required for tracking.
III Method
Figure 1: An overview of our proposed VPTracker.
Figure 2: Prompt used for Vision-Language Tracking.
III-A Problem Formulation
Visual-language tracking aims to localize an object in subsequent video frames based on the given bounding box and the corresponding language description in the first frame. Most existing trackers rely on the assumption that the target moves continuously and smoothly throughout the video. A dominant paradigm is to restrict the search region in the current frame using the target’s predicted position and size from the previous frame, allowing the model to perform localization only within a limited area. The overall pipeline can be summarized as follows:
$T=\text{Crop}(I_{1},B_{1})$ (1)
$S_{t}=\text{Crop}(I_{t},B_{t-1}),\quad t=2,\dots,N$ (2)
$B_{t}=M(T,L,S_{t})$ (3)
where $I_{t}$ denotes the $t$-th video frame; $B_{t}$ is the predicted bounding box of the target in frame $I_{t}$; $L$ is the natural language description of the target; $\text{Crop}(I_{t},B_{t-1})$ denotes the region cropped from the full image $I_{t}$ based on the previous bounding box $B_{t-1}$, after enlarging it by a certain scale; and $M(T,L,S_{t})$ denotes the localization operator that outputs the predicted bounding box $B_{t}$. This typical paradigm of cropping the search region, while avoiding excessive visual information interference and reducing the tracking difficulty for the model, also leads to target loss when encountering scenarios such as excessively fast target movement, viewpoint switching, or target occlusion.
III-B Global Tracking via MLLMs
To address the aforementioned issues, this paper proposes a global tracking method based on MLLM. Leveraging MLLM’s powerful reasoning capabilities, it searches for targets across the entire video frame given a visual template and corresponding linguistic description, and provides bounding boxes for the detected targets. As shown in Fig. 1, our method relies solely on a single multimodal large language model and does not require any additional components.
Specifically, we first crop a visual template $T$ from the initial bounding box in the first frame. Taking the second frame as an example, we draw a rectangle on $I_{2}$ based on $B_{1}$ to form a visual prompt, while the drawing details will be introduced in the next subsection. Then, the visual template $T$, the current frame with the visual prompt $I_{t}^{\prime}$, and the instruction text $L^{\prime}$ containing the language description are fed into the MLLM to obtain the tracking result.
$T=\text{Crop}(I_{1},B_{1})$ (4)
$I_{t}^{\prime}=\text{VP}(I_{t},B_{t-1}),\quad t=2,\dots,N$ (5)
$B_{t}=MLLM(T,L^{\prime},I_{t}^{\prime})$ (6)
where $\text{VP}(I_{t},B_{t-1})$ denotes drawing the visual prompt on the current frame $I_{t}$ based on $B_{t-1}$. However, compared with local tracking paradigms, the global tracking paradigm can more effectively re-acquire the target after rapid motion or temporary disappearance. Nevertheless, the enlarged search space inevitably introduces more irrelevant visual information, which increases the difficulty of accurate tracking.
III-C Visual Prompt based on Location-aware
To mitigate the performance degradation caused by visual distractions in global tracking, we introduce a location-aware visual prompting mechanism. The core idea is consistent with local tracking methods, which assume that the target’s motion trajectory is relatively smooth; thus, the target’s position in frame $t+1$ is typically close to its position in frame $t$. Unlike conventional local tracking approaches that crop a small search region—thereby discarding information outside the cropped region—we avoid restricting the model’s field of view.
Instead, we draw a rectangular box directly on the full search frame as a visual prompt, encouraging the model to prioritize searching within this region. When the target moves out of the prompted region, the model then falls back to a full-frame global search. This design preserves the advantages of global perception while effectively suppressing interference from large-scale background content.
To prevent the model from overfitting to the presence of bounding boxes, we additionally include a proportion of training samples where the visual prompt is drawn but the target lies outside the prompted region. This prevents the model from becoming overly reliant on the boxes themselves. However, during inference, we strictly draw bounding boxes based on the model’s predictions.
IV Experiment
IV-A Implementation Details
Dataset
We train our model using a mixed training set composed of TNL2K and TNLLT. The TNL2K training split contains 1,300 videos, while the TNLLT training split includes 150 long video sequences. We sample the data using a 7:3 ratio, resulting in a total of one million training samples. Each sample consists of a visual template, a search frame, and the corresponding language description.
Training and Inference
We adopted Qwen3-VL-4B as our base model. All experiments are conducted on an 8 NVIDIA H200 GPU. We use the Adamw optimizer with an initial learning rate of 2e-5 and a batch size of 128. During training, we fine-tune all layers of Qwen3-VL including vision encoder, alignment layer and llm. The entire training process takes approximately 8 hours. During inference, we adopt the default decoding and reasoning configuration of Qwen3-VL. Each frame is processed independently by feeding the full-resolution image into the model.
IV-B Comparison with State-of-the-art Methods
We test VPTracker on the large-scale test datasets of TNL2K and TNLLT, and conduct a comprehensive comparison with state-of-the-art methods..
TNLLT
The TNLLT dataset is a recently proposed benchmark for long-term vision–language tracking, designed to assess a model’s capability to localize target objects over extended video sequences using natural language descriptions. Its test set consists of 50 long and diverse video sequences. To better assess robustness, TNLLT further provides 15 challenging attributes, such as occlusion, deformation, motion blur, illumination variation, fast motion, low resolution, out-of-view, and background clutter. These rich annotations make TNLLT a comprehensive benchmark for evaluating long-term temporal modeling, and cross-modal reasoning in vision–language tracking. We compare VPTracker with existing state-of-the-art trackers in Tab.I. Compared to the state-of-the-art ReasoningTrack, which uses MLLM to update the linguistic description of the object during tracking, we achieve significant improvements over prior methods.
Figure 3: Visual comparison of tracking results with and without Visual Prompt
Figure 4: Visual comparison of tracking results with and without Visual Prompt
TABLE I: Overall tracking performance on TNLLT dataset. The best results are highlighted in bold.
Method Source Type TNLLT
SRSR PRPR NPRNPR
OSTrack [ye2022ostrack] ECCV2022 $B$ 52.1 57.3 63.6
MixFormer [cui2022mixformer] CVPR2022 $B$ 56.0 61.1 67.8
AiATrack [gao2022aiatrack] ECCV2022 $B$ 53.0 56.9 63.5
JointNLT [zhou2023joint] CVPR2023 NL 45.0 47.9 55.5
JointNLT [zhou2023joint] CVPR2023 BL 45.4 48.2 55.9
All-in-one [zhang2023all] MM2023 BL 49.7 54.8 60.6
MMTrack [zheng2023toward] TCSVT2023 BL 55.8 61.8 67.7
CiteTrack [citetracker] ICCV2023 $B$ 50.3 55.3 60.8
ROMTrack [cai2023robust] ICCV2023 $B$ 49.0 54.8 59.9
GRM [gao2023generalized] CVPR2023 $B$ 50.0 55.4 61.3
ODTrack [zheng2024odtrack] AAAI2024 $B$ 54.4 61.3 65.9
EVPTrack [shi2024evptrack] AAAI2024 $B$ 54.4 61.3 65.8
UVLTrack [ma2024unifying] AAAI2024 $B$ 53.1 58.1 64.1
AQATrack [xie2024autoregressive] CVPR2024 $B$ 55.1 61.5 66.5
UVLTrack [ma2024unifying] AAAI2024 NL 51.9 57.4 63.1
UVLTrack [ma2024unifying] AAAI2024 BL 54.4 60.3 66.2
SUTrack [sutrack] AAAI2025 BL 50.8 55.4 60.0
LMTrack [xu2025less] AAAI2025 $B$ 48.7 53.0 59.8
CTVLT [feng2025enhancing] ICASSP2025 BL 60.9 69.5 73.6
VPTracker Ours BL 62.5 72.9 73.8
TNL2K
The TNL2K dataset is a large dataset used for training and evaluating Vision-Language tracking. It contains 2000 video sequences and 663 words, of which 1300 videos are used for training and 700 videos are used for testing. It introduces two new challenges: adversarial samples and modality switches. We compare VPTracker with existing advanced trackers in Tab.II. we achieve significant improvements over prior methods.
TABLE II: Overall tracking performance on TNL2K dataset. The best results are highlighted in bold.
Method Source Type TNL2K
AUCAUC PRPR NPRNPR
TrDiMP [wang2021transformer] CVPR2021 $B$ - - -
TransT [tang2022SFTransT] CVPR2021 $B$ 50.7 51.7 57.1
AutoMatch [zhang2021learn] AAAI2021 $B$ 47.2 43.5 -
SNLT [feng2021siamese] CVPR2021 BL 27.6 41.9 -
CapsuleTNL [ma2021capsule] MM 2021 BL - - -
VLTTT [guo2022divert] NeurIPS2022 BL 53.1 53.3 -
CTRNLT [li2022cross] CVPR2022 BL 44.0 45.0 52.0
JointNLT [zhou2023joint] CVPR2023 BL 56.9 58.1 73.6
OneTrack [hong2024onetracker] CVPR2024 BL 58.0 59.1 -
QueryNLT [shao2024context] CVPR2024 BL 57.8 58.7 75.6
MAVLT [shi2025mamba] TCSVT2025 BL 63.1 66.7 -
ProVLT [10947500] TCSVT2025 BL 59.8 61.8 -
TemTrack [xie2025robust] AAAI 2025 $B$ 58.8 - -
VPTracker Ours BL 64.9 71.2 80.2
IV-C Qualitative Comparison
To better demonstrate the superiority of our method, we compared its tracking results with those of two state-of-the-art approaches: ReasoningTrack and UVLTrack. As shown in Fig. 3, our method achieves more accurate tracking performance when dealing with scenes featuring significant changes in target appearance and numerous distracting factors.
IV-D Ablation study
To validate the effectiveness of our proposed location-aware visual prompting mechanism, we conduct comparisons on the TNL2K dataset with and without the visual prompt. As shown in Tab. III, incorporating the visual prompting mechanism leads to consistent and notable improvements across metrics such as AUC, PR, and NPR. To more clearly demonstrate the effectiveness of our proposed VP mechanism, we present a visual comparison of the tracking results in Fig. 4. Since the textual description of the target may only be applicable to the first frame, global tracking without the position-aware visual prompt can lead to incorrect prediction when visually similar objects appear in the scene.
TABLE III: Overall tracking performance on TNL2K dataset.
Method TNL2K
AUCAUC PRPR NPRNPR
w. VP 64.9 71.2 80.2
w/o VP 64.3 70.1 79.2
V Conclusion
In this paper, we presented VPTracker, a global vision–language tracking framework built upon a Multimodal Large Language Model. By introducing a simple yet effective location-aware visual prompting mechanism, our method successfully injects spatial priors into the MLLM, enabling it to suppress distractors while maintaining global perception. Experiments on TNL2K and TNLLT show clear improvements in robustness and long-term stability. These results demonstrate that MLLMs, when properly guided, can serve as a strong backbone for next-generation global tracking systems.
Although VPTracker demonstrates competitive performance in vision–language tracking, it represents only an initial exploration, and we have not yet incorporated more sophisticated mechanisms to further boost its capability. We believe that MLLMs hold substantial untapped potential for the tracking community, and we hope this work encourages more researchers to explore this promising direction.

### v2

@@
-VPTracker: Global Vision-Language Tracking via Visual Prompt and MLLM
+
 Abstract
@@
 Extensive experiments show that our approach significantly enhances tracking stability and target disambiguation under challenging scenarios, opening a new avenue for integrating MLLMs into visual tracking.
-Code is available at https://github.com/jcwang0602/VPTracker.
-I Introduction
-Visual object tracking (VOT) is a fundamental problem in computer vision, supporting a wide range of downstream applications such as intelligent surveillance, human–robot interaction, augmented reality, and autonomous systems [10149530].
-Traditional visual trackers primarily rely on an initial visual template and perform frame-by-frame localization using appearance similarity [ye2022ostrack].
+Code and models are available at https://github.com/jcwang0602/VPTracker.
+keywords:
+[…]
+1 Introduction
+Visual object tracking (VOT) is a fundamental problem in computer vision, supporting a wide range of downstream applications such as intelligent surveillance, human–robot interaction, augmented reality, and autonomous systems 10149530 ; ZHAO202679 ; GE202615 ; XING202559 .
+Traditional visual trackers primarily rely on an initial visual template and perform frame-by-frame localization using appearance similarity ye2022ostrack .
 Although such methods have achieved strong short-term performance, they are inherently constrained by the limited representational capacity of a single visual template.
 As a result, their robustness deteriorates in complex scenarios involving drastic appearance changes, heavy occlusions, large camera motion, or long-term absence.
-To overcome the ambiguity of purely visual cues, vision-language tracking has recently emerged as a promising direction [11209477].
+To overcome the ambiguity of purely visual cues, vision-language tracking has recently emerged as a promising direction 11209477 .
 By incorporating natural language descriptions, trackers can leverage high-level semantic priors to better distinguish objects with similar appearance and maintain identity consistency over time.
-This paradigm has shown particular advantages in multi-object scenes and cases where visual cues alone are insufficient [wang2021towards].
+This paradigm has shown particular advantages in multi-object scenes and cases where visual cues alone are insufficient wang2021towards .
 However, existing most vision-language trackers still predominantly adopt local search strategies, restricting the search region to a small neighborhood around the previous target location.
@@
 Recent advancements in Multimodal Large Language Models (MLLMs) have opened new possibilities for rethinking the tracking pipeline.
-MLLMs exhibit remarkable capabilities in global scene understanding, cross-modal reasoning [bai2025qwen3].
+MLLMs exhibit remarkable capabilities in global scene understanding, cross-modal reasoning bai2025qwen3 .
 These strengths suggest that MLLMs could fundamentally reshape long-term tracking, allowing the model to perform global search over the entire image rather than relying on local heuristics.
@@
 This design preserves the robustness of global search while substantially reducing false positives caused by visually or semantically similar distractors.
-The experimental results on the commonly used visual language tracking datasets TNL2K [wang2021towards], and TNLLT [wang2025reasoningtrack] demonstrate that our method has achieved significant improvements in robustness, target re-identification, and long-term stability.
+The experimental results on the commonly used visual language tracking datasets TNL2K wang2021towards , and TNLLT wang2025reasoningtrack demonstrate that our method has achieved significant improvements in robustness, target re-identification, and long-term stability.
 More importantly, VPTracker introduces a new perspective of integrating MLLM into vision-language tracking tasks, indicating that future tracking will not only be driven by appearance similarity but also by advanced cross-modal reasoning.
@@
 By injecting region-level spatial priors derived from the target’s previous state, our method significantly reduces false positives and enhances target recognition while preserving global reasoning capabilities.
-  • Extensive experiments on TNL2K and TNLLT demonstrate that our approach achieves superior robustness, improved recognition capability, and strong long-term stability, establishing a new direction for integrating MLLMs into vision–language tracking..
-II Related Work
-II-A Object Tracking
-Visual object tracking (VOT) has been extensively studied over the past decade, with research evolving from correlation filter–based methods to modern deep learning–based frameworks [ye2022ostrack].
-Early trackers such as KCF [henriques2014high] and ECO [danelljan2017eco] model target appearance through handcrafted features and efficient frequency-domain operations.
-With the success of deep convolutional networks, Siamese-based trackers (e.g., SiamFC [bertinetto2016fully], SiamRPN [li2018high], SiamMask [hu2023siammask]) became dominant due to their strong template-matching capability and real-time performance.
-Transformer-based trackers (e.g., TransT [chen2021TransT], STARK [tang2022SFTransT], OSTrack [ye2022ostrack]) further improve robustness by modeling long-range dependencies with self-attention mechanisms.
+  • Extensive experiments on TNL2K and TNLLT demonstrate that our approach achieves superior robustness, improved recognition capability, and strong long-term stability, establishing a new direction for integrating MLLMs into vision–language tracking.
+2 Related Work
+2.1 Object Tracking
+Visual object tracking (VOT) has been extensively studied over the past decade, with research evolving from correlation filter–based methods to modern deep learning–based frameworks ye2022ostrack .
+Early trackers such as KCF henriques2014high and ECO danelljan2017eco model target appearance through handcrafted features and efficient frequency-domain operations.
+With the success of deep convolutional networks, Siamese-based trackers (e.g., SiamFC bertinetto2016fully , SiamRPN li2018high , SiamMask hu2023siammask ) became dominant due to their strong template-matching capability and real-time performance.
+Transformer-based trackers (e.g., TransT chen2021TransT , STARK tang2022SFTransT , OSTrack ye2022ostrack ) further improve robustness by modeling long-range dependencies with self-attention mechanisms.
 The introduction of natural language into visual tracking provides richer semantic information beyond appearance cues.
@@
 As a result, they struggle in complex environments with distractors that share similar descriptions or visual patterns.
-II-B Multimodal Large Language Models
-Multimodal Large Language Models (MLLMs), such as GPT-4V [achiam2023gpt], LLaVA [li2024llava], Qwen-VL[bai2025qwen3], and InternVL [wang2025internvl3], have exhibited strong capabilities in high-level scene understanding, cross-modal reasoning, and open-vocabulary recognition [wang2025unlocking].
+2.2 Multimodal Large Language Models
+Multimodal Large Language Models (MLLMs), such as GPT-4V achiam2023gpt , LLaVA li2024llava , Qwen-VLbai2025qwen3 , and InternVL wang2025internvl3 , have exhibited strong capabilities in high-level scene understanding, cross-modal reasoning, and open-vocabulary recognition wang2025unlocking .
 Unlike traditional vision-language models that output fixed embeddings, MLLMs are able to perform explicit reasoning, generate spatial descriptions, and interpret complex relationships between objects.
@@
 This reveals a fundamental gap between the global semantic reasoning strengths of MLLMs and the precise, temporally grounded localization required for tracking.
-III Method
+3 Method
 Figure 1: An overview of our proposed VPTracker.
 Figure 2: Prompt used for Vision-Language Tracking.
-III-A Problem Formulation
+3.1 Problem Formulation
 Visual-language tracking aims to localize an object in subsequent video frames based on the given bounding box and the corresponding language description in the first frame.
@@
 This typical paradigm of cropping the search region, while avoiding excessive visual information interference and reducing the tracking difficulty for the model, also leads to target loss when encountering scenarios such as excessively fast target movement, viewpoint switching, or target occlusion.
-III-B Global Tracking via MLLMs
+3.2 Global Tracking via MLLMs
 To address the aforementioned issues, this paper proposes a global tracking method based on MLLM.
 Leveraging MLLM’s powerful reasoning capabilities, it searches for targets across the entire video frame given a visual template and corresponding linguistic description, and provides bounding boxes for the detected targets.
-As shown in Fig.
-1, our method relies solely on a single multimodal large language model and does not require any additional components.
+As illustrated in Fig.
+1, our approach relies exclusively on a single multimodal large language model, without requiring any additional components.
+The prompt employed in our method is presented in Fig.
+2.
 Specifically, we first crop a visual template $T$ from the initial bounding box in the first frame.
@@
 Nevertheless, the enlarged search space inevitably introduces more irrelevant visual information, which increases the difficulty of accurate tracking.
-III-C Visual Prompt based on Location-aware
+3.3 Visual Prompt based on Location-aware
 To mitigate the performance degradation caused by visual distractions in global tracking, we introduce a location-aware visual prompting mechanism.
@@
 However, during inference, we strictly draw bounding boxes based on the model’s predictions.
-IV Experiment
-IV-A Implementation Details
+4 Experiment
+4.1 Implementation Details
 Dataset
@@
 Each frame is processed independently by feeding the full-resolution image into the model.
-IV-B Comparison with State-of-the-art Methods
+4.2 Comparison with State-of-the-art Methods
 We test VPTracker on the large-scale test datasets of TNL2K and TNLLT, and conduct a comprehensive comparison with state-of-the-art methods..
@@
 These rich annotations make TNLLT a comprehensive benchmark for evaluating long-term temporal modeling, and cross-modal reasoning in vision–language tracking.
-We compare VPTracker with existing state-of-the-art trackers in Tab.I.
+We compare VPTracker with existing state-of-the-art trackers in Tab.1.
 Compared to the state-of-the-art ReasoningTrack, which uses MLLM to update the linguistic description of the object during tracking, we achieve significant improvements over prior methods.
@@
 Figure 4: Visual comparison of tracking results with and without Visual Prompt
-TABLE I: Overall tracking performance on TNLLT dataset. The best results are highlighted in bold.
+Table 1: Overall tracking performance on TNLLT dataset. The best results are highlighted in bold.
 Method Source Type TNLLT
 SRSR PRPR NPRNPR
-OSTrack [ye2022ostrack] ECCV2022 $B$ 52.1 57.3 63.6
-MixFormer [cui2022mixformer] CVPR2022 $B$ 56.0 61.1 67.8
-AiATrack [gao2022aiatrack] ECCV2022 $B$ 53.0 56.9 63.5
-JointNLT [zhou2023joint] CVPR2023 NL 45.0 47.9 55.5
-JointNLT [zhou2023joint] CVPR2023 BL 45.4 48.2 55.9
-All-in-one [zhang2023all] MM2023 BL 49.7 54.8 60.6
-MMTrack [zheng2023toward] TCSVT2023 BL 55.8 61.8 67.7
-CiteTrack [citetracker] ICCV2023 $B$ 50.3 55.3 60.8
-ROMTrack [cai2023robust] ICCV2023 $B$ 49.0 54.8 59.9
-GRM [gao2023generalized] CVPR2023 $B$ 50.0 55.4 61.3
-ODTrack [zheng2024odtrack] AAAI2024 $B$ 54.4 61.3 65.9
-EVPTrack [shi2024evptrack] AAAI2024 $B$ 54.4 61.3 65.8
-UVLTrack [ma2024unifying] AAAI2024 $B$ 53.1 58.1 64.1
-AQATrack [xie2024autoregressive] CVPR2024 $B$ 55.1 61.5 66.5
-UVLTrack [ma2024unifying] AAAI2024 NL 51.9 57.4 63.1
-UVLTrack [ma2024unifying] AAAI2024 BL 54.4 60.3 66.2
-SUTrack [sutrack] AAAI2025 BL 50.8 55.4 60.0
-LMTrack [xu2025less] AAAI2025 $B$ 48.7 53.0 59.8
-CTVLT [feng2025enhancing] ICASSP2025 BL 60.9 69.5 73.6
+OSTrack ye2022ostrack ECCV22 $B$ 52.1 57.3 63.6
+MixFormer cui2022mixformer CVPR22 $B$ 56.0 61.1 67.8
+AiATrack gao2022aiatrack ECCV22 $B$ 53.0 56.9 63.5
+JointNLT zhou2023joint CVPR23 NL 45.0 47.9 55.5
+JointNLT zhou2023joint CVPR23 BL 45.4 48.2 55.9
+All-in-one zhang2023all MM23 BL 49.7 54.8 60.6
+MMTrack zheng2023toward TCSVT23 BL 55.8 61.8 67.7
+CiteTrack citetracker ICCV23 $B$ 50.3 55.3 60.8
+ROMTrack cai2023robust ICCV23 $B$ 49.0 54.8 59.9
+GRM gao2023generalized CVPR23 $B$ 50.0 55.4 61.3
+ODTrack zheng2024odtrack AAAI24 $B$ 54.4 61.3 65.9
+EVPTrack shi2024evptrack AAAI24 $B$ 54.4 61.3 65.8
+UVLTrack ma2024unifying AAAI24 $B$ 53.1 58.1 64.1
+AQATrack xie2024autoregressive CVPR24 $B$ 55.1 61.5 66.5
+UVLTrack ma2024unifying AAAI24 NL 51.9 57.4 63.1
+UVLTrack ma2024unifying AAAI24 BL 54.4 60.3 66.2
+SUTrack sutrack AAAI25 BL 50.8 55.4 60.0
+LMTrack xu2025less AAAI25 $B$ 48.7 53.0 59.8
+CTVLT feng2025enhancing ICASSP25 BL 60.9 69.5 73.6
 VPTracker Ours BL 62.5 72.9 73.8
@@
 It introduces two new challenges: adversarial samples and modality switches.
-We compare VPTracker with existing advanced trackers in Tab.II.
+We compare VPTracker with existing advanced trackers in Tab.2.
 we achieve significant improvements over prior methods.
-TABLE II: Overall tracking performance on TNL2K dataset. The best results are highlighted in bold.
+Table 2: Overall tracking performance on TNL2K dataset. The best results are highlighted in bold.
 Method Source Type TNL2K
 AUCAUC PRPR NPRNPR
-TrDiMP [wang2021transformer] CVPR2021 $B$ - - -
-TransT [tang2022SFTransT] CVPR2021 $B$ 50.7 51.7 57.1
-AutoMatch [zhang2021learn] AAAI2021 $B$ 47.2 43.5 -
-SNLT [feng2021siamese] CVPR2021 BL 27.6 41.9 -
-CapsuleTNL [ma2021capsule] MM 2021 BL - - -
-VLTTT [guo2022divert] NeurIPS2022 BL 53.1 53.3 -
-CTRNLT [li2022cross] CVPR2022 BL 44.0 45.0 52.0
-JointNLT [zhou2023joint] CVPR2023 BL 56.9 58.1 73.6
-OneTrack [hong2024onetracker] CVPR2024 BL 58.0 59.1 -
-QueryNLT [shao2024context] CVPR2024 BL 57.8 58.7 75.6
-MAVLT [shi2025mamba] TCSVT2025 BL 63.1 66.7 -
-ProVLT [10947500] TCSVT2025 BL 59.8 61.8 -
-TemTrack [xie2025robust] AAAI 2025 $B$ 58.8 - -
+TrDiMP wang2021transformer CVPR21 $B$ - - -
+TransT tang2022SFTransT CVPR21 $B$ 50.7 51.7 57.1
+AutoMatch zhang2021learn AAAI21 $B$ 47.2 43.5 -
+SNLT feng2021siamese CVPR21 BL 27.6 41.9 -
+VLTTT guo2022divert NIPS22 BL 53.1 53.3 -
+CTRNLT li2022cross CVPR22 BL 44.0 45.0 52.0
+JointNLT zhou2023joint CVPR23 BL 56.9 58.1 73.6
+OneTrack hong2024onetracker CVPR24 BL 58.0 59.1 -
+QueryNLT shao2024context CVPR24 BL 57.8 58.7 75.6
+MAVLT shi2025mamba TCSVT25 BL 63.1 66.7 -
+ProVLT 10947500 TCSVT25 BL 59.8 61.8 -
+TemTrack xie2025robust AAAI25 $B$ 58.8 - -
 VPTracker Ours BL 64.9 71.2 80.2
-IV-C Qualitative Comparison
+4.3 Qualitative Comparison
 To better demonstrate the superiority of our method, we compared its tracking results with those of two state-of-the-art approaches: ReasoningTrack and UVLTrack.
@@
 3, our method achieves more accurate tracking performance when dealing with scenes featuring significant changes in target appearance and numerous distracting factors.
-IV-D Ablation study
+4.4 Ablation study
 To validate the effectiveness of our proposed location-aware visual prompting mechanism, we conduct comparisons on the TNL2K dataset with and without the visual prompt.
 As shown in Tab.
-III, incorporating the visual prompting mechanism leads to consistent and notable improvements across metrics such as AUC, PR, and NPR.
+3, incorporating the visual prompting mechanism leads to consistent and notable improvements across metrics such as AUC, PR, and NPR.
 To more clearly demonstrate the effectiveness of our proposed VP mechanism, we present a visual comparison of the tracking results in Fig.
@@
 Since the textual description of the target may only be applicable to the first frame, global tracking without the position-aware visual prompt can lead to incorrect prediction when visually similar objects appear in the scene.
-TABLE III: Overall tracking performance on TNL2K dataset.
+Table 3: Overall tracking performance on TNL2K dataset.
 Method TNL2K
@@
 w/o VP 64.3 70.1 79.2
-V Conclusion
+5 Conclusion
 In this paper, we presented VPTracker, a global vision–language tracking framework built upon a Multimodal Large Language Model.
@@
 We believe that MLLMs hold substantial untapped potential for the tracking community, and we hope this work encourages more researchers to explore this promising direction.
+Acknowledgments
+This work was partially supported by the National Natural Science Foundation of China under Grant 62072185.